### v1

@@ 4.2 Experimental Details
      MI-Net [48]       0.6447       0.7565       0.7803       0.7933       0.7822       0.7674
      Attn-MIL [19]       0.6225       0.6213       0.7370       0.7551       0.7028       0.6342
      Gated-Attn-MIL [19]       0.6187       0.6873       0.7970       0.7831       0.6378       0.6277
      ABMIL [19]       0.6253       0.9083       0.9241       0.9237       0.8224       0.7935
      DSMIL [26]       0.4039       0.5515       0.4918       0.8258       0.6152       0.7525
      DSAGL       0.5971       0.8552       0.9037       0.9438       0.9092       0.9103
@@
      MI-Net [48]       0.6114       0.6952       0.7388       0.8640       0.9075       1.000
      Attn-MIL [19]       0.6235       0.6589       0.8832       1.000       1.000       1.000
      Gated-Attn-MIL [19]       0.5444       0.5792       0.8637       1.000       1.000       1.000
      ABMIL [19]       0.5783       0.8850       0.9955       1.000       1.000       1.000
      DSMIL [26]       0.4025       0.5174       0.5265       0.9468       0.9850       1.000
      DSAGL       0.6044       0.8203       0.9715       1.000       1.000       1.000
4.3 Comparison Methods
We compare DSAGL with a diverse set of representative multiple instance learning (MIL) approaches, encompassing classical pooling-based and attention-based models (e.g., MI-Net [48], Attention-MIL [19], ABMIL [19]), as well as more recent architectures that incorporate dual-stream mechanisms, Transformer-based global modeling, and region-aware learning strategies, including DSMIL [26], CLAM [31], TransMIL [39], IBMIL [29], and R2T-MIL [45]. For a fair comparison, both ABMIL and DSMIL adopt SimCLR [26] as their feature extractor. Experiments are conducted on the Cifar-10 and NCT-CRC datasets under varying positive instance ratios ranging from 10% to 70%. As the TCGA-Lung dataset lacks ground-truth instance-level annotations, only bag-level evaluation is performed. For several MIL baselines, part of the results are reproduced in our experimental settings, while others are directly referenced from existing studies on weakly supervised WSI classification [36].
4.4 Experimental Results
For synthetic datasets, Table 1 and Table 2 report the instance-level and bag-level classification performance of DSAGL on CIFAR-10 and NCT-CRC under varying positive patch ratios. As shown in Table 1, DSAGL exhibits strong stability across different settings, achieving instance-level AUCs of 0.9438, 0.9092, and 0.9103 at 20%, 50%, and 70%, respectively—outperforming baselines such as ABMIL and DSMIL. At the bag level, DSAGL achieves a perfect AUC of 1.000 across all three ratios, further confirming its robustness under high signal-to-noise conditions. These results demonstrate DSAGL’s ability to remain reliable in synthetic weak supervision scenarios and its resilience to noisy pseudo labels.
On the NCT-CRC dataset (Table 2), DSAGL consistently outperforms all competing methods across different positive ratios. Its instance-level AUC remains above 0.95 in all settings, reaching a peak of 0.9916 at 70%, highlighting its robustness to distributional shifts and morphological diversity. Additionally, DSAGL achieves a perfect bag-level AUC of 1.000 under all conditions, surpassing or matching the performance of state-of-the-art MIL methods. This consistent superiority confirms DSAGL’s capability to generalize across complex tissue patterns while maintaining high discriminative accuracy at both instance and bag levels.
@@ 4.4 Experimental Results
Pos. ratio 10% 20% 50% 70%
MI-Net [48] 0.6648 0.6950 0.6996 0.6899
Attn-MIL [19] 0.6621 0.6424 0.6409 0.6548
Gated-Attn-MIL [19] 0.6753 0.6333 0.6830 0.7228
ABMIL [19] 0.7410 0.8729 0.8800 0.7965
DSMIL [26] 0.3690 0.7008 0.4835 0.7399
DSAGL 0.9731 0.9538 0.9755 0.9916
@@
Pos. ratio 10% 20% 50% 70%
MI-Net [48] 0.7160 0.5185 0.7600 0.7212
Attn-MIL [19] 0.7407 0.7037 1.000 1.000
Gated-Attn-MIL [19] 0.9506 0.9506 1.000 1.000
ABMIL [19] 0.9754 1.000 0.9766 0.9894
DSMIL [26] 1.000 0.8914 0.9997 1.000
DSAGL 1.000 1.000 1.000 1.000
@@ 4.5 Ablation Study
Max-pooling [49] 0.9014
Mean-pooling [49] 0.9369
RNN-MIL [4] 0.9107
ABMIL [19] 0.9488
DSMIL [26] 0.9533
CLAM [31] 0.9607
TransMIL [39] 0.9672
IBMIL [29] 0.9689
R2T-MIL [45] 0.9701
DSAGL 0.9727
Table 4: Ablation study on the CRC-MIL dataset.
@@
       ✓        0.8021        1.000
       ✓        ✓        0.9214        1.000
       ✓        ✓        ✓        0.9348        1.000
       ✓        ✓        ✓        ✓        0.9731        1.000
Starting from a baseline model without any additional modules, the instance-level AUC is 0.7410, and the bag-level AUC is 0.9754. Introducing the dual-stream architecture improves the instance-level AUC to 0.8021, while the bag-level AUC reaches 1.000, indicating that the teacher-student collaborative optimization mechanism effectively enhances bag-level prediction performance. Incorporating the FASA further boosts the instance-level AUC to 0.9214, demonstrating the effectiveness of multi-scale saliency modeling in identifying lesion regions.
Building on this, the addition of the hybrid loss function $\mathcal{L}_{Hybrid}$ leads to a further increase in instance-level AUC to 0.9348, highlighting the robustness advantages of joint supervision through hard and soft pseudo labels in weakly supervised learning. Finally, integrating the customized feature encoder VSSMamba into the overall architecture yields the best performance, achieving an instance-level AUC of 0.9731 while maintaining a bag-level AUC of 1.000, validating the effectiveness of this module in modeling long-range dependencies and fusing multi-scale features.
@@ 4.6 Visualization and Analysis
To further interpret the classification behavior of DSAGL, we visualize the attention responses from the student branch. Specifically, Grad-CAM is applied to the final convolutional layer of the student encoder, with gradients back-propagated from the classification output to generate class-discriminative attention heatmaps. These visualizations, derived exclusively from the student network, reflect its instance-level decision preferences and provide insights into the spatial reasoning patterns learned during training.
As shown in Fig. 5, we present attention maps across four representative tissue categories in the CRC-MIL dataset. For ADI (adipose tissue), the model successfully suppresses large homogeneous regions and focuses on interstitial lines and localized boundaries, indicating robustness to non-informative areas. In MUS (muscle), attention highlights fiber boundaries and inter-fiber gaps, capturing structural cues effectively. For LYM (lymphocytes), the model accurately detects sparse lymphocyte clusters, demonstrating strong granular attention to small-scale features. In the TUM (tumor) category, attention is highly concentrated on densely packed epithelial structures with nuclear atypia and glandular distortion, reflecting the model’s capacity for complex morphological reasoning and multi-focal response.
Figure 5: Visualization of attention heatmaps on representative categories from the CRC-MIL dataset. Each pair shows the original image (left) and the corresponding attention heatmap (right) generated by the student model.
We also observe occasional failure cases. In certain MUS samples, attention mildly spreads into surrounding connective tissue, possibly due to early-stage pseudo label noise from the teacher branch. Similar issues arise in some LYM samples, suggesting that pseudo supervision consistency remains challenging in sparse lesion scenarios.
Nevertheless, the alternating training strategy and hybrid supervision in DSAGL help mitigate attention drift. By iteratively optimizing the teacher and student branches, the model achieves more stable learning and produces class-relevant attention patterns.
@@ 5 Discussion
The proposed DSAGL framework exhibits several strengths in the context of weakly supervised WSI classification. By introducing a dual-branch architecture, the model effectively narrows the supervision gap between bag-level labels and instance-level learning. The alternating training strategy facilitates bidirectional knowledge transfer between the teacher and student branches, enhancing semantic consistency and training stability. Furthermore, the integration of a Mamba-based encoder enables efficient modeling of long-range dependencies across image patches, while maintaining linear computational complexity. Experimental results across multiple datasets confirm that DSAGL achieves superior performance compared to state-of-the-art MIL methods, particularly in terms of instance-level precision and boundary localization accuracy.
Nevertheless, several limitations remain. First, in scenarios with highly imbalanced patch distributions or extremely sparse positive regions, the pseudo labels generated by the teacher branch may be noisy during early training stages, potentially affecting the convergence of the student branch. Second, the current Mamba encoder treats patch sequences as one-dimensional inputs, which may overlook crucial two-dimensional spatial structures and anatomical priors inherent in histopathological images. Third, DSAGL does not explicitly model pseudo-label uncertainty, which could reduce robustness in regions with ambiguous or noisy annotations.
Overall, DSAGL is well suited for high-resolution, structurally complex, and weakly annotated pathological images, and demonstrates strong generalization across diverse data distributions. Future work will focus on enhancing the dual-branch learning paradigm by designing more adaptive feature extractors and uncertainty-aware training objectives to further improve model robustness and interpretability.
@@ 6 Conclusion
In this work, we propose DSAGL, a novel dual-branch teacher–student framework for weakly supervised whole-slide image (WSI) classification. By integrating attention-guided pseudo labeling and an alternating training strategy, DSAGL effectively bridges the gap between coarse-grained supervision and fine-grained instance-level learning. The incorporation of a Mamba-based encoder allows efficient modeling of long-range dependencies with linear complexity, enabling the model to process large-scale instance sets while preserving global context. Additionally, the FASA module enhances the model’s ability to focus on sparse but diagnostically significant regions through multi-scale attention. Experimental results on three benchmark datasets demonstrate that DSAGL consistently outperforms existing MIL-based approaches in both instance-level and bag-level classification tasks, particularly under challenging conditions such as low positive-instance ratios and heterogeneous lesion patterns. These findings highlight the effectiveness, scalability, and generalization potential of DSAGL, making it a promising solution for real-world computational pathology applications.
@@ Acknowledgments
This work was supported by the 2024 College Students’ Innovative Entrepreneurial Training Plan Program, Program Number: 20240151.

### v2

@@ 4.4.1 Results on the CIFAR-10 Dataset
      Classical MIL Models
      MI-Net(2018) [48]       0.6447       0.7565       0.7803       0.7933       0.7822       0.7674
      RNN-MIL(2019) [4]       0.5026       0.7685       0.7466       0.7852       0.6544       0.6412
      Attention-based MIL Models
      ABMIL(2018) [19]       0.6225       0.6213       0.7370       0.7551       0.7028       0.6342
      Gated-ABMIL(2018) [19]       0.6187       0.6873       0.7970       0.7831       0.6378       0.6277
      CLAM(2021) [31]       0.6054       0.8048       0.8239       0.7771       0.6616       0.6139
      Transformer-based or Advanced MIL Models
      TransMIL(2021) [39]       0.7015       0.7360       0.8279       0.7619       0.6393       0.5938
      IBMIL(2023) [29]       0.7353       0.8532       0.8760       0.8469       0.6767       0.5167
      R2T-MIL(2024) [45]       0.6714       0.7986       0.7680       0.7945       0.6468       0.6136
      DSAGL(Ours)       0.5971       0.8552       0.9037       0.9287       0.9092       0.9103
@@
      Classical MIL Models
      MI-Net(2018) [48]       0.6114       0.6952       0.7388       0.8640       0.9075       1.0000
      RNN-MIL(2019) [4]       0.6709       0.6812       0.9230       0.9992       1.0000       1.0000
      Attention-based MIL Models
      ABMIL(2018) [19]       0.6235       0.6589       0.8832       1.0000       1.0000       1.0000
      Gated-ABMIL(2018) [19]       0.5444       0.5792       0.8637       1.0000       1.0000       1.0000
      CLAM(2021) [31]       0.4444       0.6986       0.9226       1.0000       1.0000       1.0000
      Transformer-based or Advanced MIL Models
      TransMIL(2022) [39]       0.5536       0.6952       0.9507       0.9976       1.0000       1.0000
      IBMIL(2023) [29]       0.5437       0.7633       0.8746       0.9944       1.0000       1.0000
      R2T-MIL(2024) [45]       0.6874       0.8248       0.9124       1.0000       1.0000       1.0000
      DSAGL(Ours)       0.6044       0.8303       0.9715       1.0000       1.0000       1.0000
(a) Instance-level Classification AUC.
(b) Bag-level Classification AUC.
Figure 5: Comparison of AUC for Different Classification Levels on the CIFAR-10 dataset.
(a) Instance-level Classification AUC.
(b) Bag-level Classification AUC.
Figure 6: Test AUC curves on the CIFAR-10 dataset.
4.4.2 Results on the NCT-CRC Dataset
As shown in Table 2, DSAGL consistently outperforms all competing methods across varying positive patch ratios on the NCT-CRC dataset. Its instance-level AUC remains above 0.95 in all cases, reaching as high as 0.9916 at a 70% ratio. This strong performance demonstrates DSAGL’s robustness to morphological variation and distributional shift in real-world histopathological data. At the bag level, DSAGL achieves a perfect AUC of 1.0000 under all settings, matching or exceeding the performance of the best existing MIL approaches.
Fig. 7 provides a visual comparison of instance-level and bag-level AUC trends across different ratios. Unlike competing models, which often exhibit substantial fluctuation or degradation (e.g., CLAM’s instance-level drop at 50%), DSAGL maintains consistent performance as the supervision signal changes. These results confirm the model’s strong generalization capability across heterogeneous tissue patterns and its ability to retain discriminative precision under weak supervision.
@@ 4.4.2 Results on the NCT-CRC Dataset
Positive patch ratio 10% 20% 50% 70%
Classical MIL Models
MI-Net(2018) [48] 0.6648 0.6950 0.6996 0.6899
RNN-MIL(2019) [4] 0.6148 0.6234 0.6244 0.6335
Attention-based MIL Models
ABMIL(2018) [19] 0.6621 0.6424 0.6409 0.6548
Gated-ABMIL(2018) [19] 0.6753 0.6333 0.6830 0.7228
CLAM(2021) [31] 0.6525 0.8525 0.5450 0.6446
Transformer-based or Advanced MIL Models
TransMIL(2022) [39] 0.7013 0.8530 0.6043 0.6723
IBMIL(2023) [29] 0.7532 0.8542 0.7523 0.8212
R2T-MIL(2024) [45] 0.8241 0.8720 0.7656 0.8023
DSAGL(Ours) 0.9731 0.9538 0.9755 0.9916
@@
Positive patch ratio 10% 20% 50% 70%
Classical MIL Models
MI-Net(2018) [48] 0.7160 0.5185 0.7600 0.7212
RNN-MIL(2019) [4] 0.8201 0.8324 0.9932 1.0000
Attention-based MIL Models
ABMIL(2018) [19] 0.7407 0.7037 1.0000 1.0000
Gated-ABMIL(2018) [19] 0.9506 0.9506 1.0000 1.0000
CLAM(2021) [31] 0.9828 1.0000 1.0000 1.0000
Transformer-based or Advanced MIL Models
TransMIL(2022) [39] 0.9723 1.0000 1.0000 1.0000
IBMIL(2023) [29] 1.0000 0.9978 1.0000 1.0000
R2T-MIL [45] 1.0000 1.0000 0.9997 1.0000
DSAGL(Ours) 1.0000 1.0000 1.0000 1.0000
(a) Instance-level Classification AUC.
(b) Bag-level Classification AUC.
Figure 7: Comparison of AUC for Different Classification Levels on the NCT-CRC dataset.
4.4.3 Results on the TCGA-Lung dataset
@@
Classical MIL Models
MI-Net(2018) [48] 0.9312
RNN-MIL(2019) [4] 0.9307
Attention-based MIL Models
ABMIL(2018) [19] 0.9488
Gated-ABMIL(2018) [19] 0.9588
CLAM(2021) [31] 0.9607
Transformer-based or Advanced MIL Models
TransMIL(2021) [39] 0.9672
IBMIL(2023) [29] 0.9689
R2T-MIL(2024) [45] 0.9701
DSAGL(Ours) 0.9727
4.5 Ablation Study
4.5.1 Model Architecture Design
Table 4: Ablation study on the NCT-CRC dataset.
@@ 4.5.1 Model Architecture Design
     ✓      0.8021      1.0000
     ✓      ✓      0.9214      1.0000
     ✓      ✓      ✓      0.9348      1.0000
     ✓      ✓      ✓      ✓      0.9731      1.0000
Starting from a baseline model without any additional modules, we observe an instance-level AUC of 0.7410 and a bag-level AUC of 0.9754. Introducing the dual-stream architecture improves the instance-level AUC to 0.8021, while the bag-level AUC reaches 1.0000—indicating that the teacher-student collaborative optimization mechanism effectively enhances bag-level prediction performance. Incorporating the FASA module further boosts the instance-level AUC to 0.9214, demonstrating the effectiveness of multi-scale saliency modeling in identifying diagnostically relevant lesion regions.
Building on this, the addition of the hybrid loss function $\mathcal{L}_{Hybrid}$ leads to a further increase in instance-level AUC to 0.9348, highlighting the robustness advantages of joint supervision through both hard and soft pseudo labels under weak supervision. Finally, integrating the customized feature encoder VSSMamba into the overall architecture yields the best performance, achieving an instance-level AUC of 0.9731 while maintaining a bag-level AUC of 1.0000. This result validates the effectiveness of VSSMamba in modeling long-range dependencies and integrating multi-scale features for enhanced instance-level discrimination.
4.5.2 Loss Function Design
To further investigate the contribution of each component in our loss function design, we conduct a series of ablation experiments on the CIFAR-10 dataset with a positive patch ratio of 20%, as summarized in Table 5. These experiments focus on three key factors: whether the binary cross-entropy (BCE) loss is class-balanced, whether knowledge distillation via KL divergence is applied, and what type of pseudo-label is utilized during training.
Under the full configuration (which includes weighted BCE, KL divergence, and soft pseudo-labels), our method achieves the best performance, with an instance-level AUC of 0.9438 and a bag-level AUC of 1.0000. Removing the KL divergence term leads to a noticeable drop in instance AUC to 0.9317, indicating that mimicking the teacher’s soft outputs significantly enhances the student’s representational learning.
Further eliminating the class-balancing mechanism in BCE results in a more pronounced performance degradation: the instance AUC decreases to 0.9031 and the bag-level AUC falls to 0.9814. This suggests that class imbalance adversely affects optimization, and that the reweighted BCE term contributes to better training stability.
When the soft pseudo-labels are binarized into hard labels, the instance AUC declines to 0.9123, even though KL divergence is retained. This underscores the importance of preserving uncertainty information in the pseudo-labels to guide the student more effectively. Moreover, when pseudo-labels are randomly assigned, performance drops drastically, with the instance AUC and bag-level AUC plunging to 0.6124 and 0.6749, respectively highlighting the critical role of pseudo-label quality in weakly supervised learning.
As a reference upper bound, we also evaluate a student model trained with ground-truth instance labels, which yields an instance-level AUC of 0.9420 and a bag-level AUC of 1.0000. While slightly lower than the full hybrid loss configuration, this result demonstrates that our pseudo-labeling strategy is already approaching the performance of fully supervised training.
Overall, these findings confirm that the proposed hybrid loss, which integrates confidence-aware supervision, class-balanced reweighting, and soft-label distillation, is critical for achieving robust and generalizable learning under weak supervision.
Table 5: Ablation study on the effects of pseudo-label design, KL divergence, and weighted BCE on the CIFAR-10 dataset.
Weighted BCE KL Divergence Pseudo-label Type Instance AUC Bag AUC
Soft 0.9438 1.0000
Soft 0.9317 1.0000
Soft 0.9031 0.9814
Hard 0.9123 1.0000
Random 0.6124 0.6749
GT 0.9420 1.0000
@@ 4.6 Visualization and Analysis
To further interpret the spatial reasoning behavior of DSAGL and compare its attention mechanisms with representative MIL models, we visualize class-discriminative attention responses from the student branch. Specifically, Grad-CAM is applied to the final convolutional layer of the student encoder, where gradients are backpropagated from the classification logits to generate saliency heatmaps. These visualizations highlight the regions the model attends to during inference and provide insight into its learned instance-level representations.
Fig. 8 presents a qualitative comparison of attention heatmaps across four representative tissue classes—ADI, MUS, LYM, and TUM from the NCT-CRC dataset. The top row displays the original histopathology images, followed by visualizations from four MIL models representing distinct architectural paradigms: DSAGL (ours, representing Transformer-based or advanced MIL), R2T-MIL (recent Transformer-based MIL), CLAM (attention-based MIL), and RNN-MIL (classical MIL with pooling). This selection enables a cross-generational comparison of spatial reasoning capabilities among MIL designs.
Across all tissue types, DSAGL generates attention maps that are consistently more focused, compact, and semantically aligned with diagnostic regions. In the ADI class (adipose tissue), DSAGL effectively suppresses homogeneous background and highlights interstitial boundaries, whereas other methods yield dispersed or poorly localized activation. For MUS (muscle), DSAGL sharply delineates fiber margins, while R2T-MIL and CLAM produce broader, less discriminative responses. In the LYM class (lymphocyte) , which is challenging due to its sparse distribution, DSAGL accurately localizes small lymphocyte clusters, outperforming CLAM’s overextended focus and RNN-MIL’s weak saliency. For TUM (tumor), DSAGL concentrates on epithelial abnormalities and distorted glandular structures with superior boundary precision.
Despite these strengths, some failure cases are observed. In certain MUS samples, attention maps slightly extend into adjacent connective tissue, potentially due to noisy pseudo labels in the early training stages. A similar phenomenon appears in LYM patches, suggesting pseudo-supervision inconsistencies in sparsely distributed lesions.
Overall, the alternating teacher-student optimization and hybrid supervision strategy in DSAGL effectively reduce attention drift. Through iterative pseudo-label refinement and instance-level consistency enforcement, the model achieves more stable and diagnostically meaningful attention localization. These interpretable saliency patterns further demonstrate DSAGL’s applicability and reliability in real-world weakly supervised histopathology tasks.
Figure 8: Qualitative comparison of attention heatmaps on four tissue classes (ADI, MUS, LYM, TUM) from the NCT-CRC dataset. Each row corresponds to a representative MIL model: DSAGL (Ours), R2T-MIL (Transformer-based), CLAM (Attention-based), and RNN-MIL (Classical pooling-based). DSAGL produces more localized and diagnosis-relevant attention responses across all tissue types.
@@ 5 Discussion
The proposed DSAGL framework demonstrates several strengths in the context of weakly supervised WSI classification. By introducing a dual-branch architecture, the model effectively bridges the supervision gap between bag-level labels and instance-level learning. The alternating training strategy facilitates bidirectional knowledge transfer between the teacher and student branches, enhancing semantic consistency and training stability. Furthermore, the integration of a mamba-based encoder enables efficient modeling of long-range dependencies across image patches while maintaining linear computational complexity. Experimental results across multiple datasets confirm that DSAGL outperforms state-of-the-art MIL methods, particularly in terms of instance-level precision and boundary localization accuracy.
Nevertheless, some limitations remain. First, in scenarios with highly imbalanced patch distributions or extremely sparse positive regions, the pseudo labels generated by the teacher branch may be noisy during the early stages of training, potentially affecting the convergence of the student branch. Second, the current Mamba encoder treats patch sequences as one-dimensional inputs, which may fail to capture essential two-dimensional spatial structures and anatomical priors inherent in histopathological images. Third, DSAGL does not explicitly model pseudo-label uncertainty, which could reduce robustness in regions with ambiguous or noisy annotations.
Overall, DSAGL is well suited for high-resolution, structurally complex, and weakly annotated pathological images, demonstrating strong generalization across diverse data distributions. Future work will focus on enhancing the dual-branch learning paradigm by designing more adaptive feature extractors and incorporating uncertainty-aware training objectives to improve robustness and interpretability in complex pathological contexts. In addition, extending DSAGL to multi-modal inputs or integrating it with pre-trained vision models holds promise for improving its practicality and generalizability.
@@ 6 Conclusion
In this work, we propose DSAGL, a novel dual-branch teacher-student framework for weakly supervised WSI classification. By integrating attention-guided pseudo-labeling and an alternating training strategy, DSAGL effectively bridges the gap between coarse-grained supervision and fine-grained instance-level learning. The incorporation of a Mamba-based encoder enables efficient modeling of long-range dependencies with linear complexity, allowing the model to process large-scale instance sets while preserving global context. Additionally, the FASA module enhances the model’s ability to focus on sparse yet diagnostically significant regions through multi-scale attention. Experimental results on three benchmark datasets demonstrate that DSAGL consistently outperforms existing MIL-based approaches in both instance-level and bag-level classification tasks, particularly under challenging conditions such as low positive-instance ratios and heterogeneous lesion patterns. These findings highlight the effectiveness, scalability, and generalization potential of DSAGL, making it a promising solution for real-world computational pathology applications.
@@
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgments
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Data availability
Data will be made available on request.